OpenStudy (anonymous):

Select the best possible first step to solving the system by first eliminating the y variable. 4x - 5y = 4 3x + 2y = -20 Answer Multiply the first equation by -2 and multiply the second equation by 5. Multiply the first equation by 2 and multiply the second equation by -5. Multiply the first equation by 2 and multiply the second equation by 5. None of the above

OpenStudy (anonymous):

Multiply first equation by 2 and multiply second equation by 5. 2(4x-5y=4) => 8x-10y=8 5(3x+2y=-20) =>+ 15x+10y=-100 23x = -92